OpenStudy (anonymous):

How many ways can a teacher give three different prizes to three of her 27 students?

OpenStudy (howard-wolowitz):

21 divided by 3 is seven: 7

OpenStudy (howard-wolowitz):

but

OpenStudy (howard-wolowitz):

its 9

OpenStudy (howard-wolowitz):

because 3*9=27

OpenStudy (anonymous):

How did you get 9

OpenStudy (howard-wolowitz):

Did I make that understandable

OpenStudy (anonymous):

Wait I see now thanks

OpenStudy (howard-wolowitz):

Well you take 27 and three and divide them and its 27

OpenStudy (anonymous):

Yes u did

OpenStudy (anonymous):

I don't think that's right

OpenStudy (anonymous):

\[\frac{ 27! }{ (27-3)! } = 17 550\]

OpenStudy (anonymous):

Explanation in the link.

OpenStudy (anonymous):

Thank you looking at it now

OpenStudy (anonymous):

@Howard-Wolowitz you might want to look at it too